A man bought 42 stamps, some 13¢ and some 18¢. How many of each kind did he buy if the cost was $6.66?
Found 2 solutions by addingup, macston:
Answer by addingup(3677)

Let's call the 13¢ stamps a and the 18¢ stamps b:
a+b = 42 and therefore a= 42-b (formula 1)
0.13a+0.18b= 6.66 In this formula, substitute the value of a according to formula 1:
0.13(42-b)+0.18b= 6.66 Multiply on the left to get rid of the parenthesis:
5.46-0.13b+0.18b= 6.66 Subtract 5.46 from both sides:
-0.13b+0.18b= 1.20 Add on the left:
0.05b= 1.20 Divide both sides by 0.05
b= 24 You have 24 18¢ stamps and:
42-24= 18 13¢ stamps
Check: (24 x 0.18) + (18 x 0.13)= 6.66 Correct.

Answer by macston(5194)

X=number of $0.13; Y=number of $0.18
X+Y=42
X=42-Y
$0.13X+$0.18Y=$6.66 Substitute for X.
$0.13(42-Y)+$0.18Y=$6.66
$5.46-$0.13Y+$0.18Y=$6.66 Subtract $5.46 from each side.
$0.05Y=$1.20 Divide each side by $0.05
Y=24 ANSWER1: He bought 24-18¢ stamps.
X=42-Y=42-24=18 ANSWER 2: He bought 18-13¢ stamps.
CHECK:
$0.13X+$0.18Y=$6.66
$0.13(18)+$0.18(24)=$6.66
$2.34+$4.32=$6.66
$6.66=$6.66